# Tianyang Xu, Xiaoze Liu11footnotemark: 1, Feijie Wu, Xiaoqian Wang, Jing Gao

### v1

@@ Abstract
Large Language Models (LLMs) have transformed natural language processing by learning from massive datasets, yet this rapid progress has also drawn legal scrutiny, as the ability to unintentionally generate copyrighted content has already prompted several prominent lawsuits. In this work, we introduce SUV (Selective Unlearing for Verbatim data), a selective unlearning framework designed to prevent LLM from memorizing copyrighted content while preserving its overall utility. In detail, the proposed method constructs a dataset that captures instances of copyrighted infringement cases by the targeted LLM. With the dataset, we unlearn the content from the LLM by means of Direct Preference Optimization (DPO), which replaces the verbatim copyrighted content with plausible and coherent alternatives. Since DPO may hinder the LLM’s performance in other unrelated tasks, we integrate gradient projection and Fisher information regularization to mitigate the degradation. We validate our approach using a large-scale dataset of 500 famous books (predominantly copyrighted works) and demonstrate that SUV significantly reduces verbatim memorization with negligible impact on the performance on unrelated tasks. Extensive experiments on both our dataset and public benchmarks confirm the scalability and efficacy of our approach, offering a promising solution for mitigating copyright risks in real-world LLM applications.
@@ 1 Introduction
Our approach unfolds in three primary stages. Firstly, unlike previous methods that segment text into large chunks, we employ a sliding-window mechanism to achieve fine-grained identification of memorized passages. This allows us to apply a more granular approach in distinguishing between relevant and irrelevant information within the model’s memory. Secondly, we tackle the challenge of utility degradation. Many existing approaches remove not only copyrighted content but also information that can be generalized or rephrased from the original texts. Consequently, unlearning verbatim memorization can undermine LLM utility—especially on unrelated tasks—and lead to significant performance drops as the scale of copyrighted materials increases. In contrast, our method unlearns the targeted segments via Direct Preference Optimization (DPO) (Rafailov et al., 2024), preferring a randomly generated completion over plagiarized content. Although the random completion may initially yield meaningless passages, we mitigate the degradation by integrating projected gradient techniques and Fisher information regularization to control the unlearning process and prevent dramatic LLM updates. Finally, we address dataset limitations and scalability by constructing a large-scale dataset comprising 500 famous books—predominantly copyrighted works such as Harry Potter Series and A Song of Ice and Fire Series—to faithfully mimic real-world usage. By assembling our dataset of 500 copyrighted books, our approach provides a more realistic assessment of LLM behavior under copyright constraints.
@@ 2 Related Work
Prior work has focused on examining the extent to which language models reproduce copyrighted text and how this behavior relates to the broader phenomenon of memorization. Several studies have employed various probing techniques, such as similarity measurement and masked token prediction, to evaluate verbatim reproduction and data extraction risks (Chang et al., 2023; Karamolegkou et al., 2023; D’Souza & Mimno, 2023; Hacohen et al., 2024; Nasr et al., 2023; Schwarzschild et al., 2024; Li et al., 2024; Liu et al., 2024c; Chen et al., 2024). Investigations have revealed that model capacity and training procedures play a significant role in memorization, with even state-of-the-art models outputting verbatim excerpts from their training data (Carlini et al., 2021; 2023; Biderman et al., 2023; Wei et al., 2024; Liu et al., 2024c). Concurrently, other works have introduced new evaluation metrics and defense assessments to better understand these copyright implications (Wei et al., 2024; Mueller et al., 2024; Chen et al., 2024). Further studies have explored detection methods and data curation techniques to mitigate memorization risks (Huang et al., 2024; Kim et al., 2023; Elangovan et al., 2021; Lee et al., 2022). Efforts to mitigate these issues have been broadly categorized into methods that enable models to forget specific training data associated with copyrighted content, i.e. unlearning (Liu et al., 2024b; d; Yao et al., 2023; Hans et al., 2024; Chen & Yang, 2023; Hans et al., 2024; Maini et al., 2024; Dou et al., 2025; Zhang et al., 2024), modifications at the decoding stage to steer generation away from potentially infringing text (Ippolito et al., 2023a; Xu et al., 2024b; Abad et al., 2024a), with additional enhancements explored via agents (Liu et al., 2024c). Our method falls into the unlearning category, enabling copyright compliance for open-weight LLMs. Current unlearning methods fail to handle large-scale corpus while retaining the model performance. In contrast, SUV employs targeted elimination to unlearn specific content while preserving the model’s overall knowledge and utility.
@@
Machine Unlearning in LLMs.  Machine unlearning seeks to selectively remove undesired knowledge from a pretrained model while preserving its overall capabilities (Cao & Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2022; Liu et al., 2024d; e). A central challenge is catastrophic forgetting, whereby unlearning targeted information inadvertently degrades untargeted knowledge and impairs model performance (Tarun et al., 2023; Xu et al., 2024a; Fan et al., 2024; Zhang et al., 2024). In the context of LLMs, existing mitigation strategies fall into two broad categories: parameter‑based unlearning (Jang et al., 2022; Yu et al., 2023; Zhang et al., 2023; Wu et al., 2023; Hase et al., 2023) and prompt‑based unlearning (Madaan et al., 2022; Pawelczyk et al., 2023; Thaker et al., 2024; Muresanu et al., 2024; Liu et al., 2024a). Prompt‑based methods avoid the computational expenses of retraining by steering model outputs at inference time, but they neither yield a permanently modified model nor fully guard against adversarial “jailbreak” attacks (Liu et al., 2025). In contrast, parameter‑based approaches (Zhang et al., 2024; Jia et al., 2024) produce a persistent and robust unlearned model, with the added benefit of preserving other learned knowledge. Despite these advances, it remains challenging to remove verbatim memorization of copyrighted materials in large-scale datasets without affecting performance. SUV employs a selective approach that precisely targets unwanted content while preserving learned representations, making it suitable for large-scale unlearning scenarios.
@@ 3.2 Dataset Construction
We construct a DPO dataset that meets the requirement of the above optimization problem (cf. Equation 1) . A DPO dataset consists of preference triples $(x^{(p)},y^{(w)},y^{(l)})$, where $x^{(p)}$ is the prompt, $y^{(w)}$ is the preferred continuation, and $y^{(l)}$ is the less-preferred continuation. We divide the dataset construction into two steps: (1) Plagiarism Detection, which identifies the subsequences in $\mathcal{D}_{B}$ memorized by the model, constructing $x^{(p)}$ and $y^{(l)}$ pairs, and (2) Counterfactual Generation, which generates $y^{(w)}$ as a counterfactual continuation to $x^{(p)}$.
@@
[…]
@@
For each prompt $x^{(p)}$ with its rejected continuation $y^{(l)}$, we generate a counterfactual (preferred) continuation $y^{(w)}$ using a counterfactual generation process. Specifically, we prompt the model $f_{\theta}$ with $x^{(p)}$ and instruct it to produce an alternative continuation in a different genre from $y^{(l)}$. This counterfactual continuation has both a low perplexity and a significant divergence from $y^{(l)}$. We detail the generation process in Appendix D.
@@ 3.4 Regularization Combination
In this variant, our goal is to prevent utility drop by leveraging both Fisher-based and gradient projection losses, as each plays a complementary role in protecting the parameters import to $\mathcal{D}_{U}$. Directly adding these losses is an intuitive way to maximize protection, with the gradient projection helping maintain stable update directions and the Fisher-based regularization constraining update magnitudes. However, the Fisher loss can dominate when other parts of the model are actively updated, making the optimization challenging. To overcome this, we introduce an adaptive scheduler for the Fisher regularization weight $\lambda_{\mathrm{fisher}}$. Starting from a fixed initial value, $\lambda_{\mathrm{fisher}}$ decays dynamically based on the training progress: a mild decay ($\rho_{\text{mild}}<1$) is applied if the DPO loss decreases, while a severe decay ($\rho_{\text{severe}}\ll\rho_{\text{mild}}$) is used when no improvement is observed over $R$ consecutive steps. This dynamic balancing—termed SUV-AS —ensures that both losses contribute effectively to preserving the overall utility of the model throughout training. A detailed description of the scheduler is provided in Appendix E.
@@ 4.1 Experiment settings
We evaluate SUV using two groups of benchmarks: unlearning and utility benchmarks. For assessing unlearning performance, we employ several datasets: (1) CBP-500 (Copyrighted Books (Predominantly)-500), our dataset that is comprised of texts from 500 famous books (predominantly copyrighted); (2) CBP-50, a randomly chosen subset of the CBP-500 dataset is used to ensure a fair comparison between SUV and baselines that cannot scale to 500 books; (3) CotaEval NewsQA (Wei et al., 2024), a publicly available benchmark that is comprised with texts from news articles; and (4) SSU Books (Dou et al., 2025), the 10 books used for training SSU, a baseline approach discussed below. We refer to Appendix C.1 for details on the datasets. For retaining the utility of language models, the Alpaca (Taori et al., 2023) dataset is used as the retain set of SUV, some baselines and ablation methods. For efficiency, we randomly sample 2,048 examples from Alpaca during training. To assess performance on unrelated tasks, we utilize MMLU (Hendrycks et al., 2021), a benchmark evaluating model knowledge across diverse subjects, and CSQA (Talmor et al., 2018), a benchmark with challenging multi-hop multiple-choice questions. Details can be found in Appendix C.3.
@@
SUV, baselines, and ablations are implemented with the Llama 3.1-8B model for fair comparison. More details can be found in Appendix C.4.
@@ 4.2 Experimental Results
Figure 3(a)–(c) presents the unlearning performance of our methods along with various baselines, where lower ROUGE-L scores indicate better removal of verbatim copyrighted material. We evaluate all the methods on fixed-size chunks divided from the original books (for simplicity, let’s just call them sentences) and report the number of sentences that exceed a certain ROUGE-L threshold. We analyze the unlearning performance by comparing the ratio of the sentence counts with ROUGE-L scores exceeding specific thresholds between our method and the Vanilla model. Specifically, on the CBP-500 dataset, at the 0.2 threshold, the Vanilla model registers 5,133 sentences while our method (using SUV-TV) retains only 1,176 sentences, corresponding to approximately 23% of the Vanilla count. For the 0.3 threshold, our method’s count is 169 versus 2,762 for the Vanilla model, which is only about 6% of the original. This gap further widens at the 0.5 threshold, where our method retains roughly 1.8% (23 versus 1,264 sentences) of the Vanilla count, and at the 0.7 threshold, only about 0.8% (6 versus 751 sentences) remains. These results demonstrate that our approach can selectively excise copyrighted content down to nearly 1/100 or less of the original amount when more stringent thresholds are applied. We also notice that the baselines, such as SSU and GA, tend to aggressively reduce the number of sentences above the given thresholds; however, this reduction is achieved at the expense of utility (detailed below).
@@
While some baselines may achieve low ROUGE-L counts at higher thresholds, they often do so at the cost of downstream task performance. For example, methods like SSU and GA, despite slightly lower unlearning percentages, experience significant degradation in utility metrics (e.g., CSQA scores dropping notably compared to the Vanilla model’s 0.7166). We further showcase that they have made the model unusable with a case study detailed in Section 4.2. In contrast, our SUV-AS approach consistently maintains utility scores close to the vanilla model (0.7093 for CSQA and 0.6622 for MMLU on CotaEval NewsQA), ensuring that the aggressive unlearning does not compromise overall performance. This dual-method framework—where SUV-TV excels at targeted content removal and SUV-AS preserves utility—underscores the practical advantage of our approach.
@@ 5 Conclusion
In this work, we introduced SUV, a selective unlearning framework that efficiently mitigates the verbatim memorization of copyrighted content in LLMs. By leveraging DPO together with projected gradient techniques and Fisher information regularization, SUV precisely targets and removes unwanted content while preserving the integrity of the model’s learned knowledge. Our extensive experiments on a large-scale dataset of 500 books, as well as on public benchmarks, demonstrate that SUV achieves a superior balance between effective unlearning and maintaining overall model utility. This scalability and precision make our approach a promising solution for addressing copyright risks in real-world LLM applications.
@@
In this work, our primary goal is to safeguard the intellectual property of authors and publishers against AI-driven copyright infringement. In an era of widespread information access, protecting copyrighted materials has become increasingly challenging. Our approach leverages cutting-edge technologies to identify and eliminate unauthorized reproductions of copyrighted text from generative models, while rigorously upholding the rights of content creators. To this end, we have implemented measures that ensure our method remains both respectful of intellectual property rights and anchored in principles of fairness and responsibility.

### v2

@@ Abstract
Large Language Models (LLMs) have transformed natural language processing by learning from massive datasets, yet this rapid progress has also drawn legal scrutiny, as the ability to unintentionally generate copyrighted content has already prompted several prominent lawsuits. In this work, we introduce SUV (Selective Unlearing for Verbatim data), a selective unlearning framework designed to prevent LLM from memorizing copyrighted content while preserving its overall utility. In detail, the proposed method constructs a dataset that captures instances of copyrighted infringement cases by the targeted LLM. With the dataset, we unlearn the content from the LLM by means of Direct Preference Optimization (DPO), which replaces the verbatim copyrighted content with plausible and coherent alternatives. Since DPO may hinder the LLM’s performance in other unrelated tasks, we integrate gradient projection and Fisher information regularization to mitigate the degradation. We validate our approach using a large-scale dataset of 500 famous books (predominantly copyrighted works) and demonstrate that SUV significantly reduces verbatim memorization with negligible impact on the performance on unrelated tasks. Extensive experiments on both our dataset and public benchmarks confirm the scalability and efficacy of our approach, offering a promising solution for mitigating copyright risks in real-world LLM applications. Code is available at https://github.com/xz-liu/SUV/.
@@ 1 Introduction
Our approach unfolds in three primary stages. Firstly, unlike previous methods that segment text into large chunks, we employ a sliding-window mechanism to achieve fine-grained identification of memorized passages. This allows us to apply a more granular approach in distinguishing between relevant and irrelevant information within the model’s memory. Secondly, we tackle the challenge of utility degradation. Many existing approaches remove not only copyrighted content but also information that can be generalized or rephrased from the original texts. Consequently, unlearning verbatim memorization can undermine LLM utility—especially on unrelated tasks—and lead to significant performance drops as the scale of copyrighted materials increases. In contrast, our method unlearns the targeted segments via Direct Preference Optimization (DPO) (Rafailov et al., 2024), preferring a randomly generated completion over plagiarized content. Although the random completion may initially yield meaningless passages, we mitigate the degradation by integrating projected gradient techniques and Fisher information regularization to control the unlearning process and prevent dramatic LLM updates. In contrast to additive-loss baselines such as GA and NPO that simply sum separate “forget” and “retain” objectives, our DPO formulation turns unlearning into a single, preference-based optimisation problem. This unified objective updates only the parameters most responsible for verbatim memorisation, thereby minimising collateral damage to general knowledge and downstream utility. Finally, we address dataset limitations and scalability by constructing a large-scale dataset comprising 500 famous books—predominantly copyrighted works such as Harry Potter Series and A Song of Ice and Fire Series—to faithfully mimic real-world usage. By assembling our dataset of 500 copyrighted books, our approach provides a more realistic assessment of LLM behavior under copyright constraints.
@@ 2 Related Work
Prior work has focused on examining the extent to which language models reproduce copyrighted text and how this behavior relates to the broader phenomenon of memorization. Several studies have employed various probing techniques, such as similarity measurement and masked token prediction, to evaluate verbatim reproduction and data extraction risks (Chang et al., 2023; Karamolegkou et al., 2023; D’Souza & Mimno, 2023; Hacohen et al., 2024; Nasr et al., 2023; Schwarzschild et al., 2024; Li et al., 2024; Liu et al., 2024c; Chen et al., 2024). Investigations have revealed that model capacity and training procedures play a significant role in memorization, with even state-of-the-art models outputting verbatim excerpts from their training data (Carlini et al., 2021; 2023; Biderman et al., 2023; Wei et al., 2024; Liu et al., 2024c). Concurrently, other works have introduced new evaluation metrics and defense assessments to better understand these copyright implications (Wei et al., 2024; Mueller et al., 2024; Chen et al., 2024). Further studies have explored detection methods and data curation techniques to mitigate memorization risks (Huang et al., 2024; Kim et al., 2023; Elangovan et al., 2021; Lee et al., 2022). Efforts to mitigate these issues have been broadly categorized into methods that enable models to forget specific training data associated with copyrighted content, i.e. unlearning (Liu et al., 2024b; d; Yao et al., 2023; Hans et al., 2024; Chen & Yang, 2023; Hans et al., 2024; Maini et al., 2024; Dou et al., 2025; Zhang et al., 2024), modifications at the decoding stage to steer generation away from potentially infringing text (Ippolito et al., 2023a; Xu et al., 2024b; Abad et al., 2024a), with additional enhancements explored via agents (Liu et al., 2024c) or RAG (Golatkar et al., 2024). Our method falls into the unlearning category, enabling copyright compliance for open-weight LLMs. Current unlearning methods fail to handle large-scale corpus while retaining the model performance. In contrast, SUV employs targeted elimination to unlearn specific content while preserving the model’s overall knowledge and utility.
@@
Machine Unlearning in LLMs.  Machine unlearning seeks to selectively remove undesired knowledge from a pretrained model while preserving its overall capabilities (Cao & Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2022; Liu et al., 2024d; e; Golatkar et al., 2020a; b; 2021; Dukler et al., 2023). A central challenge is catastrophic forgetting, whereby unlearning targeted information inadvertently degrades untargeted knowledge and impairs model performance (Tarun et al., 2023; Xu et al., 2024a; Fan et al., 2024; Zhang et al., 2024). In the context of LLMs, existing mitigation strategies fall into two broad categories: parameter‑based unlearning (Jang et al., 2022; Yu et al., 2023; Zhang et al., 2023; Wu et al., 2023; Hase et al., 2023) and prompt‑based unlearning (Madaan et al., 2022; Pawelczyk et al., 2023; Thaker et al., 2024; Muresanu et al., 2024; Liu et al., 2024a). Prompt‑based methods avoid the computational expenses of retraining by steering model outputs at inference time, but they neither yield a permanently modified model nor fully guard against adversarial “jailbreak” attacks (Liu et al., 2025). In contrast, parameter‑based approaches (Zhang et al., 2024; Jia et al., 2024) produce a persistent and robust unlearned model, with the added benefit of preserving other learned knowledge. Despite these advances, it remains challenging to remove verbatim memorization of copyrighted materials in large-scale datasets without affecting performance. SUV employs a selective approach that precisely targets unwanted content while preserving learned representations, making it suitable for large-scale unlearning scenarios.
@@ 3.2 Dataset Construction
We construct a DPO dataset that meets the requirement of the above optimization problem (cf. Equation 1) . A DPO dataset consists of preference triples $(x^{(p)},y^{(w)},y^{(l)})$, where $x^{(p)}$ is the prompt, $y^{(w)}$ is the preferred continuation, and $y^{(l)}$ is the less-preferred continuation. We divide the dataset construction into two steps: (1) Plagiarism Detection, which verifies the subsequences in $\mathcal{D}_{B}$ against the model’s memory, ensures $x^{(p)}$ is indeed part of $f_{\theta}$’s training set, then constructs $x^{(p)}$ and $y^{(l)}$ pairs, and (2) Counterfactual Generation, which generates $y^{(w)}$ as a counterfactual continuation to $x^{(p)}$.
@@
$\mathcal{D}_{F}=\Bigl\{(x_{k^{*}},\dots,x_{k^{*}+1})\,\Big|\,0\leq k^{*}\leq K-1,\;f_{\theta}(x_{k^{*}},\dots,x_{k^{*}+1})\geq\epsilon\Bigr\},$ (3)
@@
For each prompt $x^{(p)}$ with its rejected continuation $y^{(l)}$, we generate a counterfactual (preferred) continuation $y^{(w)}$ using a counterfactual generation process. Specifically, we prompt the model $f_{\theta}$ with $x^{(p)}$ and instruct it to produce an alternative continuation in a different genre from $y^{(l)}$. This counterfactual continuation has both a low perplexity and a significant divergence from $y^{(l)}$. We detail the generation process in Appendix F.
@@ 3.4 Regularization Combination
In this variant, our goal is to prevent utility drop by leveraging both Fisher-based and gradient projection losses, as each plays a complementary role in protecting the parameters import to $\mathcal{D}_{U}$. Directly adding these losses is an intuitive way to maximize protection, with the gradient projection helping maintain stable update directions and the Fisher-based regularization constraining update magnitudes. However, the Fisher loss can dominate when other parts of the model are actively updated, making the optimization challenging. To overcome this, we introduce an adaptive scheduler for the Fisher regularization weight $\lambda_{\mathrm{fisher}}$. Starting from a fixed initial value, $\lambda_{\mathrm{fisher}}$ decays dynamically based on the training progress: a mild decay ($\rho_{\text{mild}}<1$) is applied if the DPO loss decreases, while a severe decay ($\rho_{\text{severe}}\ll\rho_{\text{mild}}$) is used when no improvement is observed over $R$ consecutive steps. This dynamic balancing—termed SUV-AS —ensures that both losses contribute effectively to preserving the overall utility of the model throughout training. A detailed description of the scheduler is provided in Appendix G.
@@ 4.1 Experiment settings
We evaluate SUV using two groups of benchmarks: unlearning and utility benchmarks. For assessing unlearning performance, we employ several datasets: (1) CBP-500 (Copyrighted Books (Predominantly)-500), our dataset that is comprised of texts from 500 deduplicated and random famous books (predominantly copyrighted); (2) CBP-50, a randomly chosen subset of the CBP-500 dataset is used to ensure a fair comparison between SUV and baselines that cannot scale to 500 books; (3) CotaEval NewsQA (Wei et al., 2024), a publicly available benchmark that is comprised with texts from news articles; and (4) SSU Books (Dou et al., 2025), the 10 books used for training SSU, a baseline approach discussed below. We refer to Appendix E.1 for details on the datasets. For retaining the utility of language models, the Alpaca (Taori et al., 2023) dataset is used as the retain set of SUV, some baselines and ablation methods. For efficiency, we randomly sample 2,048 examples from Alpaca during training. To assess performance on unrelated tasks, we utilize MMLU (Hendrycks et al., 2021), a benchmark evaluating model knowledge across diverse subjects, and CSQA (Talmor et al., 2018), a benchmark with challenging multi-hop multiple-choice questions. Details can be found in Appendix E.3.
@@
SUV, baselines, and ablations are implemented with the Llama 3.1-8B model for fair comparison. More details can be found in Appendix E.4. Additional experiments on Qwen 3-8B and Qwen 3-14B models are also conducted to verify the effectiveness of SUV over different model architectures and sizes, as detailed in Appendix D.
@@ 4.2 Experimental Results
Figure 3(a)–(c) presents the unlearning performance of the proposed methods along with various baselines, where lower ROUGE-L scores indicate better removal of verbatim copyrighted material. We evaluate all the methods on fixed-size chunks divided from the original books (for simplicity, let’s just call them sentences) and report the number of sentences that exceed a certain ROUGE-L threshold. We analyze the unlearning performance by calculating the ratio of the sentences with ROUGE-L scores exceeding specific thresholds (the lower the better unlearning performance). The ratios obtained by the proposed method and the Vanilla model are compared. Specifically, on the CBP-500 dataset, at the 0.2 threshold, the Vanilla model registers 5,133 sentences while the proposed method (using SUV-TV) retains only 1,176 sentences, corresponding to approximately 23% of the Vanilla count. For the 0.3 threshold, the proposed method’s count is 169 (versus 2,762 for the Vanilla model), which is only about 6% of the original. This gap further widens at the 0.5 threshold, where the proposed method retains roughly 1.8% (23 versus 1,264 sentences) of the Vanilla count, and at the 0.7 threshold, only about 0.8% (6 versus 751 sentences) remains. These results demonstrate that the proposed approach can selectively remove copyrighted content down to nearly 1/100 or less of the original amount when more stringent thresholds are applied. We also notice that the baselines, such as SSU and GA, tend to aggressively reduce the number of sentences above the given thresholds; however, this reduction is achieved at the expense of utility (detailed below).
@@
While some baselines may achieve low ROUGE-L counts at higher thresholds, they often do so at the cost of downstream task performance degradation. For example, methods like SSU and GA, despite having slightly lower unlearning metrics, experience significant degradation in utility metrics. For example, their CSQA scores drop notably compared to the Vanilla model’s 0.7166. We further show that they have made the model unusable with a case study detailed in Section 4.2. In contrast, our SUV-AS approach consistently maintains utility scores close to the vanilla model (0.7093 for CSQA and 0.6622 for MMLU on CotaEval NewsQA), ensuring that the aggressive unlearning does not compromise overall performance. This dual-method framework—where SUV-TV excels at targeted content removal and SUV-AS preserves utility—underscores the practical advantage of the proposed approach.
@@ 5 Conclusion
In this work, we introduced SUV, a selective unlearning framework that efficiently mitigates the verbatim memorization of copyrighted content in LLMs. By leveraging DPO together with projected gradient techniques and Fisher information regularization, SUV precisely targets and removes unwanted content while preserving the integrity of the model’s learned knowledge. Our extensive experiments on a large-scale dataset of 500 books, as well as on public benchmarks, demonstrate that SUV achieves a superior balance between effective unlearning and maintaining overall model utility. This scalability and precision make the proposed approach a promising solution for addressing copyright risks in real-world LLM applications.
@@
In this work, our primary goal is to safeguard the intellectual property of authors and publishers against AI-driven copyright infringement. In an era of widespread information access, protecting copyrighted materials has become increasingly challenging. The proposed approach leverages cutting-edge technologies to identify and eliminate unauthorized reproductions of copyrighted text from generative models, while rigorously upholding the rights of content creators. To this end, we have implemented measures that ensure the proposed method remains both respectful of intellectual property rights and anchored in principles of fairness and responsibility.
@@
Acknowledgements
This work was funded in part by the National Science Foundation under Grant No. IIS-2141037.